OpenStudy (anonymous):

Select the system of equations that corresponds to the given graph. graph of two lines. Both lines have a negative slope and go through the points 2 comma 0 and 0 comma 4 4x + y = 6 6x + 3y = 6 −3x + y = 2 4x + 2y = 4 2x + y = 4 6x + 3y = 12 4x + 2y = 6 6x + 3y = 12

OpenStudy (anonymous):

there is only one system where both lines go through the point (0, 4): 2x+y=4 6x+3y=12 notice that if you solve either line for y (to get into slope intercept form) you get the same line (if the lines share two points, then they are the same line) 2x+y=4 y=-2x+4 also 6x+3y=12 3y=-6x+12 y=-2x+4 the third system is the answer

OpenStudy (mikezack123):

@mandolino i had this same question, I chose the third system and i got it wrong :(